If the distribution is really N(5.43, 0.54), what proportion of observations would be less than 5.05?
Use Table A and give your answer to 4 decimal places.

thanks

Answer by Boreal(15235)

z value of interest is (x-mean)/sd
That is z=(5.05-5.43)/0.54=(-0.38)/(0.54)=-0.7037
probability is 0.2408